# Low Complexity Channel Estimation for RIS-Assisted THz Systems with Beam Split

### v1

@@ III-D CBS-GAMP Based Full CSI Estimation
IV Low-Complexity Channel Estimation by BSAD
The CBS-GAMP algorithm obtained from Section III-B ensures the accuracy of channel estimation. However, this algorithm may introduce additional computational complexity due to its execution at each SC. It is obvious that the frequency-dependent spatial directions affected by beam split effect share a common physical support. Inspired by this observation, we propose a novel BSAD scheme to reduce computational complexity. In our proposed algorithm, we only employ the CBS-GAMP to acquire the full CSI within a small subset of SCs, namely the first and the $M/2$-th SCs [^1] The selection of SC indices is demonstrated in Section IV-B. Then, employing the obtained full CSI of the subset of SCs, we design EnM algorithm and DS-MUSIC algorithm to determine the common support at BS and RIS, respectively. Finally, employing the obtained common support and the frequency-dependent nature introduced by beam split effect, we transform the channel estimation problem at the remaining SCs into a simple linear regression problem, which can be easily solved using LS method.
IV-A EnM Based DoDs Estimation at BS
In this subsection, we derive the DoDs at BS from the full CSI $\hat{\mathbf{h}}_{k}^{\mathrm{cas}}[\bar{m}],\bar{m}\in\left\{1,M/2\right\}$ estimated by the CBS-GAMP algorithm. Recalling (60), the estimated full CSI is consisted of two parts, i.e., the total CBS dictionary $\mathbf{\Pi}_{\bar{m}}$ and the angle-excluded vector $\mathbf{x}_{k,\bar{m}}$, which are known separately. Based on the fact that the $q$-th ($q=1,\dots,Q_{\mathrm{tot}}$) entry in $\mathbf{x}_{k,\bar{m}}$ corresponds to the $q$-th column in $\mathbf{\Pi}_{\bar{m}}(:,q)$, the common support can be easily determined. However, the inevitable noise and estimation errors will cause power leakage in the estimated $\hat{\mathbf{x}}_{k,\bar{m}}$, leading to more than $\Sigma_{k}$ non-zero entries. Therefore, we propose a novel EnM method to determine the common support $\zeta^{\mathrm{BS}}$ and then calculate the corresponding physical DoDs at BS, as summarized in Algorithm 2.
Algorithm 2 EnM Based DoDs Estimation at BS
0:  Angle-excluded matrix $\mathbf{x}_{k,\bar{m}},\bar{m}\in\left\{1,M/2\right\}$, the path number $L$ of BS-RIS channel and the quantization level of CBS dictionaries $Q_{\mathrm{R}}$ and $Q_{\mathrm{T}}$.
@@ IV-A EnM Based DoDs Estimation at BS
6:  $\zeta^{\mathrm{BS}}=\left\{q|\mathop{\mathrm{maxk}}\limits_{q}\left(\boldsymbol{\epsilon},L\right)\right\}$
7:  $\hat{\boldsymbol{\varphi}}=\bar{\boldsymbol{\varphi}}\left(\zeta^{\mathrm{BS}}\right)$.
7:  $\zeta^{\mathrm{BS}}$ and $\left\{\hat{\varphi_{l}}\right\}_{l=1}^{L}$.
In the proposed EnM method, we first establish the normalized energy basis vector $\tilde{\boldsymbol{\epsilon}}_{\bar{m}}$, in which each entry is the normalized results of the corresponding element in $\mathbf{x}_{k,\bar{m}}$ (Step $2$). As mentioned in Section III-A, $\mathbf{\Pi}_{\bar{m}}$ can be regarded as a block matrix of $Q_{\mathrm{T}}$ blocks, with each block contains $Q_{\mathrm{B}}$ columns. Since each block in $\mathbf{\Pi}_{\bar{m}}$ is aligned with a physical DoD at BS, $\tilde{\boldsymbol{\epsilon}}_{\bar{m}}$ can be partitioned in the same way. By accumulating the $\ell_{2}$-norm for each block of the two normalized energy basis vectors, the merged-energy vector $\boldsymbol{\epsilon}$ can establish a one-to-one correspondence with the physical DoDs at BS in the entire beam space (Step $3-5$). Subsequently, the indices corresponding to the $L$ largest elements in $\boldsymbol{\epsilon}$ can constitute the common support $\zeta^{\mathrm{BS}}$ (Step $6$). Finally, with knowledge of $\zeta^{\mathrm{BS}}$, the DoDs $\left\{\hat{\varphi_{l}}\right\}_{l=1}^{L}$ is easily obtained according to the sampling rule of the CBS dictionary (Step $7$). Fig. 3 depicts an example of the estimated energy spectrum employing the proposed EnM method. Assuming that the cascaded channel contains $L=3$ BS-RIS paths with the common support $\zeta^{\mathrm{BS}}=\left\{3,4,24\right\}$ and SNR $=20$ dB. It can be observed from the spectrum that the energy is distributed surrounding the estimated DoDs. Furthermore, the energy is concentrated around the actual direction and can be easily distinguished, which validates the effectiveness of the proposed EnM method.
@@
IV-B DS-MUSIC Based Coupled DoAs/DoDs Estimation at RIS
The proposed EnM method ensures the acquisition of the DoDs at BS. However, due to the coupling between the DoAs and DoDs at RIS, the aforementioned method can be challenging to obtain angular information at RIS. Therefore, in this subsection we propose a novel DS-MUSIC algorithm to acquire the coupled DoAs/DoDs at RIS.
@@ IV-B DS-MUSIC Based Coupled DoAs/DoDs Estimation at RIS
Firstly, with knowledge of the estimated DoDs $\left\{\hat{\varphi_{l}}\right\}_{l=1}^{L}$ from Section IV-A, the estimated cascaded channel $\hat{\mathbf{H}}_{k}^{\mathrm{cas}}[\bar{m}]$ can be projected onto the subspace of ARM of BS using the pseudo-inverse, which is given by
@@
$\begin{split}\mathbf{U}_{k,\bar{m}}&\triangleq\hat{\mathbf{H}}_{k}^{\mathrm{cas}}[\bar{m}](\hat{\mathbf{A}}_{N_{\mathrm{T}},\bar{m}}^{\mathsf{H}}\left(\boldsymbol{\varphi}\right))^{\dagger}\\ &=\mathrm{Diag}\left(\mathbf{h}_{k}[\bar{m}]\right)\mathbf{A}_{\mathrm{R},\bar{m}}\left(\boldsymbol{\psi}\right)\mathbf{\Upsilon}_{\bar{m}},\end{split}$ (61)
@@
where $\mathbf{A}_{N}^{\dagger}\mathbf{A}_{N}=\mathbf{I}_{L}$ and $\left(\mathbf{A}_{N}^{\mathsf{H}}\right)^{\dagger}=(\mathbf{A}_{N}^{\dagger})^{\mathsf{H}}$. Besides, the rest-CSI matrix $\mathbf{U}_{k,\bar{m}}\in\mathbb{C}^{N_{\mathrm{R}}\times L}$ can be block-partitioned column-wise as $\mathbf{U}_{k,\bar{m}}=\left[\mathbf{u}_{k,\bar{m},1},\cdots,\mathbf{u}_{k,\bar{m},L}\right]$. By substituting (7) into (61), the $l$-th column $\mathbf{u}_{k,\bar{m},l}$ can be expressed as
@@
$\begin{split}\mathbf{u}_{k,\bar{m},l}&=\!\upsilon_{\bar{m},l}\left(\mathbf{A}_{\mathrm{R},\bar{m}}^{*}\left(\boldsymbol{\vartheta}_{k}\right)\boldsymbol{\beta}_{k,\bar{m}}\right)\bullet\mathbf{a}_{\mathrm{R},\bar{m}}\left(\psi_{l}\right)\\ &=\!\upsilon_{\bar{m},l}\!\left[\mathbf{a}_{\mathrm{R},\bar{m}}\left(\psi_{l}\!-\!\vartheta_{k,1}\!\right)\!,\!\cdots\!,\mathbf{a}_{\mathrm{R},\bar{m}}\left(\psi_{l}\!-\!\vartheta_{k,J_{k}}\!\right)\right]\!\boldsymbol{\beta}_{k,\bar{m}},\end{split}$ (62)
@@
where $\upsilon_{l}$ is the $\left(l,l\right)$-th entry in $\mathbf{\Upsilon}_{m}$. Define $\mathbf{A}_{\mathrm{R},\bar{m}}\left(\boldsymbol{\gamma}_{k,l}\right)\triangleq\left[\mathbf{a}\left(\gamma_{k,l,1}\right),\cdots,\mathbf{a}\left(\gamma_{k,l,J_{k}}\right)\right]\in\mathbb{C}^{N_{\mathrm{R}}\times J_{k}}$ as the coupled ARM where $\gamma_{k,l,j}=\psi_{l}-\theta_{k,j}\in\left[-2,2\right],l=1,\dots,L,j=1,\dots,J_{k}$ and $\boldsymbol{\chi}_{k,\bar{m},l}\triangleq\upsilon_{\bar{m},l}\boldsymbol{\beta}_{k,\bar{m}}\in\mathbb{C}^{J_{k}}$ as the amplitude vector, (62) is rewritten as
$\mathbf{u}_{k,\bar{m},l}=\mathbf{A}_{\mathrm{R},\bar{m}}\left(\boldsymbol{\gamma}_{k,l}\right)\boldsymbol{\chi}_{k,\bar{m},l}.$ (63)
@@
To find the coupled physical directions $\left\{{\hat{\gamma}}_{k,l,j}\right\}_{j=1}^{J_{k}}$ in (63), this issue can be regarded as a frequency estimation finding problem. Inspired by this, we propose a novel DS-MUSIC algorithm as described in Algorithm 3, where $\mathrm{peaks}\left(\boldsymbol{x},k\right)$, $\mathrm{maxk}\left(\boldsymbol{x},k\right)$ and $\mathrm{mink}\left(\boldsymbol{x},k\right)$ return $k$ largest peaks, $k$ largest elements and $k$ smallest elements of vector $\boldsymbol{x}$, respectively.
@@
11:        $\mathbf{p}_{k,\bar{m},l}^{\mathrm{MUSIC}}\!\!=\!\log\!\left[\!\min\!\left(\tilde{\mathbf{p}}_{k,\bar{m},l}^{\mathrm{MUSIC}}\!\right)\!\!,\tilde{\mathbf{p}}_{k,\bar{m},l}^{\mathrm{MUSIC}}\!\!\!,\min\!\left(\tilde{\mathbf{p}}_{k,\bar{m},l}^{\mathrm{MUSIC}}\!\right)\!\right]$.
@@
16:  $\zeta^{\mathrm{RIS}}=\mathcal{Q}_{1}\cup\cdots\cup\mathcal{Q}_{L}$.
17:  $\hat{\boldsymbol{\gamma}}_{k}=\bar{\boldsymbol{\gamma}}\left(\zeta^{\mathrm{RIS}}\right)$.
17:  $\zeta^{\mathrm{RIS}}$ and $\hat{\boldsymbol{\gamma}}_{k}$.
Due to the non-strict orthogonality of $\boldsymbol{\chi}_{k,\bar{m},l}$ and limited measurements at each SC, some energy from physical directions will leak into the noise space, making it challenging to distinguish physical directions. To address this issue, we preprocess the projected signal $\mathbf{u}_{k,\bar{m},l}$ by dividing the RIS array into multiple overlapping subarrays, with each subarray contains $N_{\mathrm{sub}}$ reflecting elements. There are $N_{\mathrm{a}}=N_{\mathrm{R}}-N_{\mathrm{sub}}+1$ subarrays in total which satisfies $N_{\mathrm{a}}>N_{\mathrm{sub}}$ and $N_{\mathrm{a}},N_{\mathrm{sub}}>J_{k}$. The rest-CSI $\mathbf{u}_{k,\bar{m},l,n}$ processed by the $n$-th subarray can be expressed as
where $\mathbf{u}_{k,\bar{m},l,n}\triangleq\mathbf{u}_{k,\bar{m},l}\left(n:n+N_{\mathrm{sub}}-1\right)$ and $\mathbf{A}_{\mathrm{R},\bar{m},n}\triangleq\mathbf{A}_{\mathrm{R},\bar{m}}\left(n:n+N_{\mathrm{sub}}-1\right)$. The covariance matrix $\mathbf{R}_{k,\bar{m},l,n}^{\mathrm{sub}}\in\mathbb{C}^{N_{\mathrm{sub}}\times N_{\mathrm{sub}}}$ is given by
The covariance matrix $\mathbf{R}_{k,\bar{m},l}^{u}\in\mathbb{C}^{N_{\mathrm{sub}}\times N_{\mathrm{sub}}}$ of $\mathbf{u}_{k,\bar{m},l}$ is calculated by averaging all $N_{\mathrm{a}}$ $\mathbf{R}_{k,\bar{m},l,n}^{\mathrm{sub}}$, which is given by
@@
where $\tilde{\mathbf{U}}_{k,\bar{m},l}\triangleq\left[\mathbf{u}_{k,\bar{m},l,1},\dots,\mathbf{u}_{k,\bar{m},l,N_{\mathrm{a}}}\right]\in\mathbb{C}^{N_{\mathrm{sub}}\times N_{\mathrm{a}}}$ is the concatenation of the subspace supported vectors $\mathbf{u}_{k,\bar{m},l,n}$ in (64). The eigendecomposition of $\mathbf{R}_{k,\bar{m},l}^{u}$ yields
@@
where $\tilde{\boldsymbol{\Omega}}_{k,\bar{m},l}\triangleq\left[\tilde{\boldsymbol{\omega}}_{k,\bar{m},l,1},\cdots,\tilde{\boldsymbol{\omega}}_{k,\bar{m},l,N_{\mathrm{sub}}}\right]\in\mathbb{C}^{N_{\mathrm{sub}}\times N_{\mathrm{sub}}}$ is the eigenbasis matrix which contains $N_{\mathrm{sub}}$ eigenvectors and $\tilde{\boldsymbol{\lambda}}_{k,\bar{m},l}\in\mathbb{C}^{N_{\mathrm{sub}}}$ is the eigenvalue vector. It is obvious that $\mathbf{R}_{k,\bar{m},l}^{u}$ is a Hermitian matrix of all its $N_{\mathrm{sub}}$ eigenvectors are orthogonal to each other. Arrange the eigenvalues of $\mathbf{R}_{k,\bar{m},l}^{u}$ in descending order and organize the corresponding eigenvectors accordingly, it can be rewritten as
@@
where $\mathbf{\Omega}_{k,\bar{m},l}$ and $\boldsymbol{\lambda}_{k,\bar{m},l}$ are the sorted eigenbasis matrix and eigenvalue vector, respectively. Subsequently, $\mathbf{\Omega}_{k,\bar{m},l}\triangleq[\mathbf{\Omega}_{k,\bar{m},l}^{\mathrm{S}}\,\,\mathbf{\Omega}_{k,\bar{m},l}^{\mathrm{N}}]$ can be divided into two parts with $\mathbf{\Omega}_{k,\bar{m},l}^{\mathrm{S}}\in\mathbb{C}^{N_{\mathrm{sub}}\times J_{k}}$ and $\mathbf{\Omega}_{k,\bar{m},l}^{\mathrm{N}}\in\mathbb{C}^{N_{\mathrm{sub}}\times\left(N_{\mathrm{sub}}-J_{k}\right)}$. Wherein, $J_{k}$ largest eigenvalues corresponding to the directions of largest variability span the signal subspace $\mathbf{\Omega}_{k,\bar{m},l}^{\mathrm{S}}$, while the remaining $N_{\mathrm{sub}}-J_{k}$ eigenvectors span the noise subspace $\mathbf{\Omega}_{k,\bar{m},l}^{\mathrm{N}}$, which is orthogonal to the signal subspace. Define the sub-ARV $\mathbf{a}_{\mathrm{R},\bar{m}}^{\mathrm{sub}}(\gamma)\in\mathbb{C}^{N_{\mathrm{sub}}}$ as the first $N_{\mathrm{sub}}$ row of $\mathbf{a}_{\mathrm{R},\bar{m}}(\gamma)$. Similar to the definition of sub-ARV, we define $\mathbf{\Xi}_{\mathrm{R},\bar{m}}^{\mathrm{sub}}\in\mathbb{C}^{N_{\mathrm{sub}}\times Q_{\mathrm{B}}}$ as the sub-CBS dictionary which consists of the first $N_{\mathrm{sub}}$ rows. By employing the sub-CBS dictionary $\mathbf{\Xi}_{\mathrm{R},\bar{m}}^{\mathrm{sub}}$, one can scan the entire beam space and calculate the MUSIC-pseudo-spectrum (PS) for the beams aligned with any sampled direction. Define $\mathbf{\tilde{p}}_{k,\bar{m},l,q}^{\mathrm{MUSIC}}\triangleq[\tilde{p}_{k,\bar{m},l,1}^{\mathrm{MUSIC}},\dots,\tilde{p}_{k,\bar{m},l,Q_{\mathrm{B}}}^{\mathrm{MUSIC}}]^{\mathsf{T}}$ as the MUSIC-PS, wherein $\tilde{p}_{k,\bar{m},l,q}^{\mathrm{MUSIC}}$ is the spectral value, which follows a $\ell_{2}$-norm form as

### v2

@@ III-C Common Angle Estimation
This method implies that we could perform the estimation for each subcarrier independently to obtain the channel estimates for all subcarriers. However, this requires implementing the estimation algorithm $M$ times, which would result in an unacceptably high computational complexity. While the beam split effect complicates the channel estimation process, it also offers some insights. It is noticed in (11) that different subcarriers share common physical directions, with the only different term $\eta_{m}$. This indicates that the physical angles remain constant across different subcarriers. Inspired by this, we perform the cascaded channel estimation on only a small subset of subcarriers to acquire the common physical angles. We select $\hat{\mathbf{h}}_{k}^{\mathrm{cas}}[m]$ for $m\in\{1,M/2\}$. Given this estimation, we can further estimate the common directions, i.e., the DoDs at the BS and the coupled directions at the RIS.
III-C1 Common DoDs Estimation at BS
Given the estimated cascaded channel $\hat{\mathbf{h}}_{k}^{\mathrm{cas}}[m]=\mathbf{\Pi}_{m}\hat{\mathbf{x}}_{k,m},\bar{m}\in\left\{1,M/2\right\}$, it is noticed that the $q$-th ($q=1,\dots,Q_{\mathrm{tot}}$) entry in $\hat{\mathbf{x}}_{k,\bar{m}}$ corresponds to the $q$-th column in $\mathbf{\Pi}_{\bar{m}}(:,q)$. This means that given the accurate estimation of $\hat{\mathbf{x}}$, we only need to determine its non-zero positions corresponding to the columns in the matrix $\mathbf{\Pi}_{\bar{m}}$. Since $\mathbf{\Pi}_{\bar{m}}\triangleq\left(\mathbf{C}_{\mathrm{T},\bar{m}}^{*}\otimes\mathbf{\Xi}_{\mathrm{R},\bar{m}}\right)$ is in one-to-one correspondence with the ARVs, this relationship allows us to directly obtain the DoDs at BS. However, inevitable noise and estimation errors can cause power leakage in the estimated $\hat{\mathbf{x}}_{k,\bar{m}}$, which will lead to more than $\Sigma_{k}$ non-zero entries. To address this, we propose a heuristic method to determine the accurate non-zero positions, referred to as common physical support, and calculate the corresponding physical DoDs at the BS, as summarized in Algorithm 1.
Algorithm 1 DoDs Estimation at BS
0:  Angle-excluded matrix $\mathbf{x}_{k,\bar{m}},\bar{m}\in\left\{1,M/2\right\}$, the path number $L$ of BS-RIS channel and the quantization level of overcomplete dictionaries $Q_{\mathrm{R}}$ and $Q_{\mathrm{T}}$.
@@ III-C1 Common DoDs Estimation at BS
6:  $\mathcal{I}^{\mathrm{BS}}=\left\{q|\mathop{\mathrm{maxk}}\limits_{q}\left(\boldsymbol{\epsilon},L\right)\right\}$
7:  $\hat{\boldsymbol{\varphi}}=\bar{\boldsymbol{\varphi}}\left(\mathcal{I}^{\mathrm{BS}}\right)$.
7:  $\mathcal{I}^{\mathrm{BS}}$ and $\left\{\hat{\varphi_{l}}\right\}_{l=1}^{L}$.
In the proposed method, we first establish the normalized energy basis vector $\tilde{\boldsymbol{\epsilon}}_{\bar{m}}$, where each entry is the normalized value of the corresponding element in $\hat{\mathbf{x}}_{k,\bar{m}}$ (Step $2$). As observed in Section III-B3, $\mathbf{\Pi}_{\bar{m}}$ can be regarded as a block matrix of $Q_{\mathrm{T}}$ blocks, each containing $Q_{\mathrm{B}}$ columns. Since each block in $\mathbf{\Pi}_{\bar{m}}$ is aligned with a physical DoD at BS, $\tilde{\boldsymbol{\epsilon}}_{\bar{m}}$ can be partitioned in the same way. By accumulating the $\ell_{2}$-norm for each block, we form the merged-energy vector $\boldsymbol{\epsilon}$ to establish a one-to-one correspondence with the physical DoDs at BS (Step $3-5$). The indices corresponding to the $L$ largest elements in $\boldsymbol{\epsilon}$ will form the common physical support $\mathcal{I}^{\mathrm{BS}}$ (Step $6$). Finally, the DoDs $\left\{\hat{\varphi_{l}}\right\}_{l=1}^{L}$ can be obtained using the knowledge of $\mathcal{I}^{\mathrm{BS}}$ and the sampling rule of the CBS dictionary (Step $7$).
Fig. 4 depicts an example of the estimated energy spectrum employing the proposed method. The cascaded channel is assumed to contain $L=3$ BS-RIS paths with the common physical support $\mathcal{I}^{\mathrm{BS}}=\left\{3,4,24\right\}$ and SNR $=20$ dB. The spectrum reveals that the energy is distributed surrounding the estimated DoDs which closely matches the actual directions and can be easily distinguished. This validates the effectiveness of the proposed method.
@@
III-C2 Common Coupled Angle Estimation at RIS
@@
With the estimated DoDs $\left\{\hat{\varphi_{l}}\right\}_{l=1}^{L}$, the estimated cascaded channel $\hat{\mathbf{H}}_{k}^{\mathrm{cas}}[\bar{m}]$ can be projected onto the subspace of ARM of BS using the pseudo-inverse, which can be given by
@@
$\begin{split}\mathbf{U}_{k,\bar{m}}&\triangleq\hat{\mathbf{H}}_{k}^{\mathrm{cas}}[\bar{m}](\hat{\mathbf{A}}_{N_{\mathrm{T}},\bar{m}}^{\mathsf{H}}\left(\boldsymbol{\varphi}\right))^{\dagger}\\ &=\mathrm{Diag}\left(\mathbf{h}_{k}[\bar{m}]\right)\mathbf{A}_{\mathrm{R},\bar{m}}\left(\boldsymbol{\psi}\right)\mathbf{\Sigma}_{\bar{m}},\end{split}$ (26)
@@
where $\mathbf{A}_{N}^{\dagger}\mathbf{A}_{N}=\mathbf{I}_{L}$, and $\mathbf{U}_{k,\bar{m}}\triangleq\left[\mathbf{u}_{k,\bar{m},1},\cdots,\mathbf{u}_{k,\bar{m},L}\right]\in\mathbb{C}^{N_{\mathrm{R}}\times L}$ denotes the rest-CSI matrix, which is only related to the AoAs and AoDs at RIS. By substituting (7) into (26), the $l$-th column $\mathbf{u}_{k,\bar{m},l}$ can be expressed as
@@
$\begin{split}\mathbf{u}_{k,\bar{m},l}&=\Sigma_{\bar{m},l}\left(\mathbf{A}_{\mathrm{R},\bar{m}}^{*}\left(\boldsymbol{\vartheta}_{k}\right)\boldsymbol{\beta}_{k,\bar{m}}\right)\bullet\mathbf{a}_{\mathrm{R},\bar{m}}\left(\psi_{l}\right)\\ &=\Sigma_{\bar{m},l}\left[\mathbf{a}_{\mathrm{R},\bar{m}}\left(\gamma_{k,l,1}\!\right),\dots,\mathbf{a}_{\mathrm{R},\bar{m}}\left(\gamma_{k,l,J_{k}}\!\right)\right]\!\boldsymbol{\beta}_{k,\bar{m}},\end{split}$ (27)
where $\Sigma_{\bar{m},l}$ denotes the $\left(l,l\right)$-th element in $\Sigma_{\bar{m}}$, and $\gamma_{k,l,j}=\psi_{l}-\theta_{k,j}\in\left[-2,2\right]$ represents the coupled angle for $l=1,\dots,L$ and $j=1,\dots,J_{k}$.
@@
By defining $\mathbf{A}_{\mathrm{R},\bar{m}}\left(\boldsymbol{\gamma}_{k,l}\right)\triangleq\left[\mathbf{a}\left(\gamma_{k,l,1}\right),\cdots,\mathbf{a}\left(\gamma_{k,l,J_{k}}\right)\right]\in\mathbb{C}^{N_{\mathrm{R}}\times J_{k}}$ as the coupled ARM, (27) can be rewritten as
$\mathbf{u}_{k,\bar{m},l}=\mathbf{A}_{\mathrm{R},\bar{m}}\left(\boldsymbol{\gamma}_{k,l,j}\right)\boldsymbol{\chi}_{k,\bar{m},l},$ (28)
where $\boldsymbol{\chi}_{k,\bar{m},l}\triangleq\Sigma_{\bar{m},l}\boldsymbol{\beta}_{k,\bar{m}}\in\mathbb{C}^{J_{k}}$ denotes the amplitude vector with the path gains and phase shifts for each coupled angle.
The expression in (28) represents a typical DoA estimation problem, which can be effectively addressed using the MUSIC algorithm due to its high resolution and robustness against noise [29]. However, there are two key challenges in directly applying the MUSIC algorithm. Firstly, the effective angle range for $\boldsymbol{\gamma}_{k,l,j}$, which is the difference between two angles $\psi_{l}$ and $\theta_{k,j}$, is doubled compared to the classical angle range. Moreover, the relative frequency parameter $\eta_{m}$ introduced by beam split effect further expands this range. This results in an increased number of potential estimated angles, making it difficult to distinguish the true physical directions. Secondly, the non-strict orthogonality of $\boldsymbol{\chi}_{k,\bar{m},l}$ and limited measurements at each SC can cause energy leakage into the noise space, which further complicates the identification of physical directions.
@@
To address these challenges, we combine the information from two selected subcarriers and apply spatial smoothing preprocessing to the projected rest-CSI matrix $\mathbf{U}_{k,\bar{m}}$. Inspired by the smooth-MUSIC algorithm, we propose a novel DS-MUSIC algorithm to accurately estimate the coupled angles, as described in Algorithm 2, where $\mathrm{peaks}\left(\boldsymbol{x},k\right)$, $\mathrm{maxk}\left(\boldsymbol{x},k\right)$ and $\mathrm{mink}\left(\boldsymbol{x},k\right)$ return $k$ largest peaks, $k$ largest elements and $k$ smallest elements of vector $\boldsymbol{x}$, respectively.
We first preprocess the projected signal $\mathbf{u}_{k,\bar{m},l}$ by dividing the RIS array into multiple overlapping subarrays, with each subarray containing $N_{\mathrm{sub}}$ reflecting elements. There are $N_{\mathrm{a}}=N_{\mathrm{R}}-N_{\mathrm{sub}}+1$ subarrays in total, which satisfies $N_{\mathrm{a}}>N_{\mathrm{sub}}$ and $N_{\mathrm{a}}$, $N_{\mathrm{sub}}>J_{k}$. The rest-CSI $\mathbf{u}_{k,\bar{m},l,n}$ processed by the $n$-th subarray can be expressed as
where $\mathbf{u}_{k,\bar{m},l,n}\triangleq\mathbf{u}_{k,\bar{m},l}\left(n:n+N_{\mathrm{sub}}-1\right)$ and $\mathbf{A}_{\mathrm{R},\bar{m},n}\triangleq\mathbf{A}_{\mathrm{R},\bar{m}}\left(:,n:n+N_{\mathrm{sub}}-1\right)$. Therefore, the covariance matrix $\mathbf{R}_{k,\bar{m},l,n}^{\mathrm{sub}}\in\mathbb{C}^{N_{\mathrm{sub}}\times N_{\mathrm{sub}}}$ can be written as
@@
11:        $\mathbf{p}_{k,\bar{m},l}=\left[\min\left(\tilde{\mathbf{p}}_{k,\bar{m},l}\right),\tilde{\mathbf{p}}_{k,\bar{m},l},\min\left(\tilde{\mathbf{p}}_{k,\bar{m},l}\right)\right]$.
@@
16:  $\mathcal{I}^{\mathrm{RIS}}=\mathcal{Q}_{1}\cup\cdots\cup\mathcal{Q}_{L}$.
17:  $\hat{\boldsymbol{\gamma}}_{k}=\bar{\boldsymbol{\gamma}}\left(\mathcal{I}^{\mathrm{RIS}}\right)$.
17:  $\mathcal{I}^{\mathrm{RIS}}$ and $\hat{\boldsymbol{\gamma}}_{k}$.
To implement the spatial smoothing, the covariance matrix $\mathbf{R}_{k,\bar{m},l}^{u}\in\mathbb{C}^{N_{\mathrm{sub}}\times N_{\mathrm{sub}}}$ of $\mathbf{u}_{k,\bar{m},l}$ is calculated by averaging all $N_{\mathrm{a}}$ subarray covariance matrices $\mathbf{R}_{k,\bar{m},l,n}^{\mathrm{sub}}$, which can be expressed by
@@
where $\tilde{\mathbf{U}}_{k,\bar{m},l}\triangleq\left[\mathbf{u}_{k,\bar{m},l,1},\dots,\mathbf{u}_{k,\bar{m},l,N_{\mathrm{a}}}\right]\in\mathbb{C}^{N_{\mathrm{sub}}\times N_{\mathrm{a}}}$ is the concatenation of the subspace-supported vectors $\mathbf{u}_{k,\bar{m},l,n}$ as defined in (29). Then, the eigendecomposition of $\mathbf{R}_{k,\bar{m},l}^{u}$ can be expressed as
@@
where $\tilde{\boldsymbol{\Omega}}_{k,\bar{m},l}\triangleq\left[\tilde{\boldsymbol{\omega}}_{k,\bar{m},l,1},\cdots,\tilde{\boldsymbol{\omega}}_{k,\bar{m},l,N_{\mathrm{sub}}}\right]\in\mathbb{C}^{N_{\mathrm{sub}}\times N_{\mathrm{sub}}}$ denotes the eigenbasis matrix which contains $N_{\mathrm{sub}}$ eigenvectors and $\tilde{\boldsymbol{\lambda}}_{k,\bar{m},l}\in\mathbb{C}^{N_{\mathrm{sub}}}$ is the eigenvalue vector. Note that $\mathbf{R}_{k,\bar{m},l}^{u}$ is a Hermitian matrix, which means that all its $N_{\mathrm{sub}}$ eigenvectors are orthogonal to each other. By arranging the eigenvalues of $\mathbf{R}_{k,\bar{m},l}^{u}$ in descending order and organizing the corresponding eigenvectors accordingly, it can be rewritten as
@@
where $\mathbf{\Omega}_{k,\bar{m},l}$ and $\boldsymbol{\lambda}_{k,\bar{m},l}$ denote the sorted eigenbasis matrix and eigenvalue vector, respectively. Therefore, the matrix $\mathbf{\Omega}_{k,\bar{m},l}\triangleq[\mathbf{\Omega}_{k,\bar{m},l}^{\mathrm{S}}\,\,\mathbf{\Omega}_{k,\bar{m},l}^{\mathrm{N}}]$ can be divided into two parts: $\mathbf{\Omega}_{k,\bar{m},l}^{\mathrm{S}}\in\mathbb{C}^{N_{\mathrm{sub}}\times J_{k}}$ and $\mathbf{\Omega}_{k,\bar{m},l}^{\mathrm{N}}\in\mathbb{C}^{N_{\mathrm{sub}}\times\left(N_{\mathrm{sub}}-J_{k}\right)}$. The part $\mathbf{\Omega}_{k,\bar{m},l}^{\mathrm{S}}$ consists of the $J_{k}$ largest eigenvalues corresponding to the directions of largest variability, which spans the signal subspace. The remaining $N_{\mathrm{sub}}-J_{k}$ eigenvectors form $\mathbf{\Omega}_{k,\bar{m},l}^{\mathrm{N}}$, spans the noise subspace and is orthogonal to the signal subspace.
To calculate the MUSIC pseudo-spectrum, the practical implementation requires sampling the estimated angles. We define $\mathbf{\tilde{p}}_{k,\bar{m},l,q}\triangleq[\tilde{p}_{k,\bar{m},l,1},\dots,\tilde{p}_{k,\bar{m},l,Q_{\mathrm{B}}}]^{\mathsf{T}}$ as the MUSIC pseudo-spectrum, where $\tilde{p}_{k,\bar{m},l,q}$ is the sampling spectral value for $q=1,\dots,Q_{\mathrm{B}}$. To align these sampled values with our proposed algorithm, we adjust the angle sampling within the range of the CBS dictionary. Therefore, we define the sub-CBS dictionary $\mathbf{\Xi}_{\mathrm{R},\bar{m}}^{\mathrm{sub}}\in\mathbb{C}^{N_{\mathrm{sub}}\times Q_{\mathrm{B}}}$ as the first $N_{\mathrm{sub}}$ rows of the original CBS dictionary in (15).According to (17), each element in the sub-CBS dictionary $\xi_{\mathrm{R},\bar{m},q}^{\mathrm{sub}}$ can be derived as
$\xi_{\mathrm{R},\bar{m},q}^{\mathrm{sub}}\triangleq e^{-j2\pi\eta_{m}\left(\frac{q}{Q_{\mathrm{R}}}-1\right)\mathbf{n}_{\mathrm{sub}}},q=1,\dots,Q_{\mathrm{B}},$ (34)
where $\mathbf{n}_{\mathrm{sub}}=\left[1,\cdots,N_{\mathrm{sub}}\right]$.
By mapping the sampled beams $\xi_{\mathrm{R},\bar{m},q}^{\mathrm{sub}}$ onto the noise space $\mathbf{\Omega}_{k,\bar{m},l}^{\mathrm{N}}$, the MUSIC algorithm constructs the pseudo-spectrum by evaluating the inverse of the projection’s norm at each potential direction, which can be given as